The lengths of stay (in days) are recorded for patients admitted with coded diagnosis for COPD. They are as follows: 13, 11, 3, 12, 2, 10, 17, 24, 20, 1, 35, 5, 5, 3, 31, 30, 39, 1, 2, 6, 1, 1, 10, 2, 21, 11, 5, 3, 1, 2, 17, 6, 8, 5, 1, 2, 4, 4, 6, 19, 2, 5, 34, 2, 4, 2, 6, 9, 61, 40, 26, 21, 14, 8, 4, 13, 17, 51, 15, 56, 6, 2, 33, 14, 3, 40, 24, 3, 5, 8, 18, 5, 31, 8, 21, 14, 5, 26, 10, 6, 57, 51, 2, 22, 8, 6.

Determine: Percentage of scores that fall between 1 through 8 days. And the percentage of scores that fall between 40 through 61 days.
Answer #1

There are total of 86 observations

Number of scores that fall between 1 through 8 days = 45

So, Percentage of scores that fall between 1 through 8 days = (45/86)*100 = 52.32558 %

Number of scores that fall between 40 through 61 days = 7

So, the percentage of scores that fall between 40 through 61 days = (7/86)*100 = 8.139535 %